40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Mar 19 23:44:00 2018, gautam, Update, IOO, IMC loop checkup IMC_RF_noise.pdfIMC_OLTF_20180320.pdf
    Reply  Tue Mar 20 16:53:03 2018, gautam, Update, IOO, IMC loop checkup IMC_RF_noise.pdf
       Reply  Tue Mar 20 21:08:03 2018, gautam, Update, IOO, IMC loop checkup 
          Reply  Tue Mar 20 22:44:45 2018, gautam, Update, IOO, IMC loop checkup 
             Reply  Wed Mar 21 15:52:45 2018, gautam, Update, IOO, MC error point calibration IMC_PDHdisc_20180321.pdfIMC_PDHdisc785_20180321.pdfIMC_oscope.pdf
                Reply  Wed Mar 21 17:31:21 2018, gautam, Update, IOO, MC error point calibration IMC_RF_noise_calib.pdf
                   Reply  Wed Mar 21 21:13:44 2018, gautam, Update, IOO, IMC noise budget IMC_RF_noise_calib.pdf
                      Reply  Tue Mar 27 01:39:44 2018, gautam, Update, IOO, PSL noise eater was off IR_ALS_20180326.pdf
                         Reply  Tue Mar 27 19:32:03 2018, arijit, Update, IOO, PSL noise eater was off MCREFL_TF.pdfMCREFL_SPECTRUM.pdf
                            Reply  Tue Mar 27 23:37:35 2018, gautam, Update, IOO, MC REFL PD removed IMG_6955.JPG
                               Reply  Wed Mar 28 21:31:39 2018, gautam, Update, IOO, MC REFL PD removed 
                            Reply  Wed Mar 28 21:47:37 2018, arijit, Update, IOO, MCREFL_PD Optical response measurement LISO_fit_with_info.pdfLISO_fit.zip
                               Reply  Thu Mar 29 17:57:36 2018, arijit, Update, IOO, MCREFL_PD Optical response measurement NOISE_EATER_On_OFF.pdf
                                  Reply  Fri Mar 30 06:14:31 2018, rana, Update, IOO, MCREFL_PD Optical response measurement 
                                     Reply  Fri Mar 30 22:37:36 2018, Kevin, Update, IOO, MCREFL_PD Optical response measurement REFLPD_DemodBoard.pdf
                      Reply  Thu Apr 5 04:36:56 2018, Kevin, Update, IOO, Coil driver noise MC3.pdfMC2.pdfCoilDriver.pdf
                         Reply  Thu Apr 5 12:13:18 2018, Koji, Update, IOO, Coil driver noise 
                            Reply  Fri Apr 6 21:39:09 2018, gautam, Update, IOO, MC2 suspension health checkup MC2_Freeswinging.pdf
                               Reply  Mon Apr 9 18:46:03 2018, gautam, Update, IOO, Further debugging MC_Freeswinging.pdfIMC_NB_20180409.pdf
                                  Reply  Tue Apr 10 14:28:44 2018, gautam, Update, IOO, Further debugging 
                                     Reply  Wed Apr 11 01:34:31 2018, gautam, Update, IOO, Activities today 
                                        Reply  Thu Apr 12 18:12:49 2018, gautam, Update, ALS, NPRO channels hijacked IR_ALS_20180412.pdfBeatMouthDrift.pngETMX_20180416.jpg
                                           Reply  Fri Apr 13 11:02:41 2018, gautam, Update, ALS, EX fiber polarization drift EX_pol_drift.png
                                              Reply  Fri Apr 13 17:56:26 2018, gautam, Update, ALS, Fibers switched out 
                                                 Reply  Sat Apr 14 14:42:09 2018, gautam, Update, ALS, Fibers switched out PSLdrift.png
                                                    Reply  Tue Apr 17 14:08:29 2018, gautam, Update, ALS, Fibers switched out PSL_fluctuations.png
                                                       Reply  Fri Apr 20 00:26:34 2018, gautam, Update, ALS, Fibers switched out 
                                                          Reply  Sat Apr 21 20:25:12 2018, gautam, Update, ALS, PSL fiber pickoff status PSL-beatMouthPickoff.png
                                  Reply  Thu Apr 19 00:03:51 2018, gautam, Update, IOO, More IMC NBing IMC_NB_20180409.pdfIMC_coils_20180418.pdf
                                     Reply  Thu Apr 19 17:15:35 2018, gautam, Update, IOO, More IMC NBing IMC_NB_20180419.pdfIMC_coilDriverNoises_20180419.pdf
                                        Reply  Sun May 20 23:43:50 2018, gautam, Update, IOO, Coil driver noise re-measurement MC1_coilDriver.pdfMC2_coilDriver.pdfMC3_coilDriver.pdfMC_coilDriverNoises.tgzActuationChainNoiseMeas.pdf
                                           Reply  Tue May 22 17:26:25 2018, gautam, Update, IOO, MC1 Coil Driver pulled out 
                                              Reply  Tue May 22 23:28:01 2018, gautam, Update, IOO, MC1 Coil Driver pulled out 
                                                 Reply  Wed May 23 17:58:48 2018, gautam, Update, IOO, MC1 Coil Driver pulled out MC1_monitorTF.pngMC1_ULnoise.pdf
                                                    Reply  Wed May 30 10:17:46 2018, gautam, Update, IOO, MC1 Coil Driver pulled out EM_pickup.pdfcoilDriverNoiseComparison.pdf
                            Reply  Fri Apr 6 22:23:53 2018, Kevin, Update, IOO, Coil driver noise MC2_coil_driver.pdf
Message ID: 13696     Entry time: Wed Mar 21 15:52:45 2018     In reply to: 13694     Reply to this: 13697
Author: gautam 
Type: Update 
Category: IOO 
Subject: MC error point calibration 

As discussed at the meeting, I decided to calibrate the MC error point into physical units of Hz/rtHz (a.k.a. the PDH discriminant). This is to facilitate the debugging of the hypothesized excess IMC sensing noise. I did this as follows.

  1. Trust the POX calibration that was last updated in Aug 2017.
  2. Hook up spare DAC channel (piped from LSC rack to 1X2) to IN2 of IMC CM board.
  3. Inject excitation into MC error point via "IN2" input of the common board. For an excitation of 30cts with the IN2 gain at -32dB, I was able to see a peak in the calibrated X arm control signal that was ~x10 above the nominal noise level around 150Hz without seeing any nonlinear coupling effects in the DTT spectrum (I'm assuming 150Hz is sufficiently above the UGF of the X arm locking loop such that no loop correction is necessary).
  4. Took a spectrum of the IMC error signal, teed off into the SR785 at the I output of the demod board with the same linewidth as the DTT spectrum.
  5. Confirmed that without any excitation,
  6. Did the math to make these two peaks line up. The resulting calibration is: 13kHz/Vrms.

Math details:

  • DTT peak height @ 150 Hz with Hanning window, 25 avgs = 1.97e-4 nm/rtHz (See Attachment #1).
  • X arm cavity length = 37.79m, using which the above number becomes 1.47 Hz/rtHz.
  • Peak height in SR785 spectrum with Hanning window, 25 avgs = 1.13e-4 Vrms/rtHz (See Attachment #2).
  • Dividing, we get 13kHz/Vrms.

Using this, I can now make up a noise budget of sorts for the IMC sensing.


gautam 20180327 4.30pm: I re-checked the PDH error signal calibration using the oscilloscope method. Attachment #3 shows the PDH I and Q error signals and also the output of the RF monitor port, during a TEM00 flash. This attachment should be compared to Attachment #2 of elog 12822, and the answer lines up quite well. From my Finesse model of the IMC, I calculated that the x-axis of the PDH horn-to-horn is ~12.3kHz. Comparing to the top row of Attachment #3, I get a PDH error signal calibration of ~12.4kHz/Vrms, which lines up well with the number quoted above. So I trust my calibration, and hence, the y-axis of my noise budgets in reply to this elog.

Attachment 1: IMC_PDHdisc_20180321.pdf  3 kB  Uploaded Wed Mar 21 17:45:18 2018  | Hide | Hide all
IMC_PDHdisc_20180321.pdf
Attachment 2: IMC_PDHdisc785_20180321.pdf  177 kB  Uploaded Wed Mar 21 17:49:48 2018  | Hide | Hide all
IMC_PDHdisc785_20180321.pdf
Attachment 3: IMC_oscope.pdf  285 kB  Uploaded Tue Mar 27 17:40:40 2018  | Hide | Hide all
IMC_oscope.pdf
ELOG V3.1.3-